# What Does Neuro Mean to Cardio? Investigating the Role of Clinical Specialty Data in Medical LLMs

### v2

@@ 3.2 Results
Looking at the results of models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (55.4%) was achieved by the model trained on Infect. If the improvements were due to knowledge injection, we would expect to see best performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other pretrained LLMs in §A.12 and note that similar conclusions hold whereas with different transfer patterns across specialties.
@@ 4 Discussion
Where do improvements come from?
We hypothesize that the performance improvements in Table 2 are primarily due to domain shifting from general LLMs to the clinical domain, rather than due to the injection of new clinical knowledge.
To assess that, we analyze token probability changes for clinical terms linked to a single clinical specialty between a baseline model and the same model fine-tuned on data from each different specialty. In Fig. 2 we show the average log-probabilities of medical terms across specialties as predicted by Mistral-v0.2 as the base model and its fine-tuned variants for each of the six specialties. Regardless of the specialty used for fine-tuning, we observe increased token log-probabilities for clinical terms specific to the fine-tuning specialty as well as terms associated with other specialties.
This aligns with our observations in the cross-specialty evaluation (§3.2) that, regardless of the specialty on which the model is trained and tested, consistent improvements are achieved over the base model. We note that token log-probabilities for terms from different specialties differ in range, likely reflecting the pre-trained model’s existing knowledge distribution. This also seems to support our hypothesis that fine-tuning shifts the domain rather than injects new domain knowledge.
We further fine-tune Mistral-v0.2 on 3 Social Sciences subsets of MMLU (public relations, security studies, and sociology). There remains a gap between the average log probabilities of all six specialty terms generated by the model compared to the model trained on S-MedQA, ensuring that the observed probability increase is a result of domain-specific adaptation rather than a mere consequence of additional training (See §A.13 for detail).
Figure 2: Negative log-probabilities for clinically relevant tokens between baseline Mistral-v0.2 and the same model further fine-tuned on each specialty data. Each group represents tokens categorized into different clinical specialties. Each color means that the same model is further fine-tuned on each specialty data.
@@ 5 Conclusions
In this paper, we introduce S-MedQA, the first medical question-answering dataset annotated across 15 distinct clinical specialties. Using S-MedQA, we demonstrate that fine-tuning with medical QA data enhances LLM performance, with the improvements primarily attributed to domain shifting rather than knowledge injection. However, the precise impact of different types of QA data remains unclear (e.g., complexity of the QA pair), and we recommend further research to investigate the role of fine-tuning in the medical domain.
@@
We limited our experiments to the medical domain. However, the findings’ generalizability to other knowledge-intense domains is unknown. Further research is needed to investigate the role of fine-tuning and instruction data in other domains.
@@ A.2 Prompts used for specialty classification
A.3 Distribution of predicted specialties
Figure 3: The distribution of all specialties classified by GPT-3.5. The dark blue specialties are the 15 we finally included in our benchmark.
In Figure 3 we show the distribution of samples across specialties. We show the 15 specialties we include in S-MedQA in dark blue, comprising in total 70.0% / 70.7% / 70.1% of the entire train / validation / test sets. We do not include the rest of the specialties due to too few samples.
@@ A.3 Distribution of predicted specialties
A.5 Methods and evaluation of multi-specialty categorization
We curate a high-quality calibration set of 300 samples, annotated independently by a medical expert (300 samples) and three medical students (100 samples each). We define the ground truth as the union of the expert and the students’ annotations to minimize the risk of missing potentially relevant specialties.
For each example in the calibration set, we compute log-probabilities for all 55 candidate specialties using GPT-3.5. This is done iteratively: at each step, we extract the top-k (with k=20) next-token predictions from the model, and extend the input with tokens that could lead to valid specialty names. This process continue until either all 55 specialties were generated — allowing us to collect their corresponding token-level log-probabilities — or the next token required for a specialty does not appear in the top-20 predictions, in which case we assign that specialty a probability of zero. For each successfully generated specialty, we compute a score as the average of its next-token log-probabilities, and normalize these scores to obtain a probability distribution over the 55 specialties. We then define the conformal score as $1-P$, where $P$ is the normalized probability of a given specialty, and use this to construct prediction sets with a 95% confidence level ($\alpha$=0.05). Since each question may have multiple correct specialties, we select $P$ as the highest predicted probability among the set of annotated correct specialties for evaluation.
We report key metrics computed on the calibration dataset to evaluate the performance of our multi-specialty categorization framework. The observed coverage rate, which measures the proportion of samples where the true labels are included in the predicted label set, was 95%, indicating a good calibration. The average number of predicted labels per sample was $2.46$, with an F-1 score of $0.60$.
A.6 Clinical specialty benchmark description
In Table 3, we show the 15 specialties we include in S-MedQA, as well as the respective numbers of samples in their train sets. The quantities in each column represent the number of samples obtained via majority voting with 3+, 4+, and 5 prompts.
Number of Votes (out of 5) 3+ 4+ 5
Cardiology 2928 2571 2090
Gastroenterology 1893 1590 1176
Obstetrics and gynecology 10671 10590 10457
Neurology 2700 2249 1495
Infectious diseases 1857 1397 845
Pediatrics 8566 8365 8181
Emergency medicine 1164 787 451
Hematology 1882 1642 1264
Endocrinology 2204 1972 1443
Nephrology 1501 1302 970
Respiratory medicine 1201 799 376
Rheumatology 1000 862 624
Dermatology 536 444 328
Psychiatry 532 466 374
Orthopedics 372 292 217
Total 40007 35328 30309
Table 3: Train sets description. Number of samples of the 15 specialties using different minimum numbers of votes (3+, 4+, 5) in the train sets included in S-MedQA.
A.7 Accuracy vs. coverage trade-off of GPT-3.5 single specialty predictions
In Table 4, we list the results of our manual validation. The accuracy when using only a single prompt ranges from 73% to 80%. We also report the coverage and accuracy after applying different majority voting strategies, (i.e., at least 3, 4, or 5 responses agreed upon). Only the questions that obtain at least this number of votes are kept. We note that there is an inherent trade-off between accuracy and coverage when deciding the threshold to use for majority voting.
In practice, we release all individual prompt predictions, as well as three versions of the dataset for majority voting with a minimum of 3+, 4+, and 5 votes. A coverage of $89.1$% of the data leads to clinical specialties that are $90.8$% accurate, whereas in the other side of the spectrum, we can obtain an accuracy of $97.8$% while the coverage decreases to $49.2$%. By sharing multiple versions of S-MedQA, we cater to different users’ needs. Users can then use more data (coverage of $89.1$%) if their use-case can cope with mistakes in the order of $10$% (majority voting $3+$); if the use-case requires data akin to gold-standard, i.e., error-free, users can use majority voting $5$ (which basically requires all 5 prompts to agree for an example to be included), which provides an accuracy of $97.8$%.
Prompts
#1 #2 #3 #4 #5
Accuracy(%) 76.0 72.8 73.0 73.8 80.2
Table 4: Accuracy of each prompt. Prompts #$1$ to #$5$ are shown in Figures 59 in §A.2.
A.8 Term overlap analysis
Test Sets Cardio(%) Gastro(%) Infect(%) Neuro(%) Obstetrics(%) Pediatrics(%)
Train Sets Cardio 67.5 31.4 38.3 31.0 25.7 31.5
Gastro 30.8 64.3 43.4 32.1 24.8 36.7
Infect 26.1 30.9 64.1 27.5 28.4 40.8
Neuro 27.5 27.4 41 62.3 23.5 33.8
Obstetrics 25.1 26.9 56.5 27.2 62.9 40.8
Pediatrics 32.8 33.4 44.8 28.8 34.7 59.2
Table 5: Medical term overlap ratios between train/test sets across specialties and within the same specialty (in percentage). Each column represents a test set as the denominator for overlap ratio calculations, while each row represents the test sets used for matching overlapping terms. Underlined scores indicate overlaps between the training and test sets within the same specialty, while other scores represent overlaps between different specialties.
To perform an overlap analysis of medical terms, we first extract clinically relevant terms in S-MedQA, map these terms onto all of their relevant clinical specialties within the scope of the top 6 specialties we selected for further experiments. We leverage SNOMED-CT, [^6] https://www.snomed.org/ a medical ontology system with all medical concepts, and Human Phenotype Ontology (HPO; Castellanos et al., 2024), which provides a standardized vocabulary of phenotypic abnormalities encountered in human disease. We include 4 types of medical concepts in SNOMED-CT: disorder, finding, procedure, and observable entity.
The high-level SNOMED-CT concepts we use for each speciality is: Disorder of cardiovascular system (disorder), Disorder of digestive system (disorder), Infectious disease (disorder), Disorder of female reproductive system (disorder), Disorder of fetus and/or mother during labor (disorder), Disorder of nervous system (disorder), Behavioral and emotional disorder with onset in childhood (disorder), Disorder of fetus and/or newborn (disorder), Developmental disorder (disorder).
We directly map disorders onto one or more specialties by traversing through their ancestors until they match any high-level specialty-specific concepts as above. We found no simple mechanism to map other types of terms (i.e., findings, procedures, observable entities) to specialties. We thus first search these terms in HPO to identify their relevant disorders, and then used these disorders similarly to map onto clinical specialties.
The average overlap in train/test splits across specialties (e.g., Cardio’s train vs. Neuro’s test) is 32.8%, indicating that different specialties only share limited common specialty-specific terminology, hence there will be minor “knowledge leakage” when evaluating cross-specialty performances. Meanwhile, the average overlap between a same speciality’s train/test sets is 63.4%. This shows that the training sets contains sufficient knowledge coverage to solve the questions in the test sets.
A.9 Prompt for LLM tuning and inferring
A. Atenolol
B. Diltiazem
C. Propafenone
D. Digoxin
Answer: [/INST] D. Digoxin
A.10 Training settings and hyperparameters
Figure 4: The illustration of first token probability, string matching, and our approach (classifier) to evaluating LLMs performance on S-MedQA. We use text output instead of first token probability for evaluation because first token probability suffers heavily from selection bias in multiple-choice question answering (Wang et al., 2024). However, string matching does not work in some cases. Our classifier trained on Mistral-v0.2 works successfully with an accuracy of 96.5%.
We use LoRA (Hu et al., 2021) on all projection layers for the fine-tuning process in all experiments. The hyperparameters are as follows: learning rate=2e-5, rank=32, alpha=16, dropout rate=0.1, batch size=8.
A.11 LLM evaluation protocol
In all test sets, we shuffle the answers 5 times for each sample and add all these 5 entries to the final test set in case the model prefers an option due to its position (Zheng et al., 2023). To further improve the reliability of results, we follow Wang et al. (2024) to generate the entire answer with the model and train a classifier to match model outputs to the options in a post-hoc step, instead of using the maximum probability of options {A, B, C, D} with a single next-token prediction step. More concretely, we randomly select 150 training samples and generate answers for these with all four LLMs (Llama2-7b, Llama2-13b, Mistral-v0.1, and Mistral-v0.2), resulting in $600$ responses. We manually annotate all the responses with the right options and use these annotations to train a Mistral-Instruct-v0.2 model as the classifier, with $400$ ($200$) train (test) samples. Our classifier achieves 96.5% accuracy and we use it in all experiments.
A.12 Cross-specialty evaluation results for Mistral-7B-v0.1, Llama-2-7B and -13B, Llama-3.1-8B, Llama-3.2-3B, Bio-Medical-LLaMA-3-8B, and OLMo-2-1124-7B-Instruct
@@ A.11 LLM evaluation protocol
In Tables 78,  9,  10,  11,  12,  13 we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, Mistral-7b-instruct-v0.1, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Bio-Medical-LLaMA-3-8B, and and OLMo-2-1124-7B-Instruct in addition to our main results (in §3.2). Here we also observe that the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty for Llama models.
A.13 Average token log probabilities gap between model fine-tuned on Social Science subsets vs. S-MedQA
We fine-tune Mistral-v0.2 on three Social Sciences subsets of MMLU—public relations, security studies, and sociology—which are entirely unrelated to the medical domain. We compare the average token log-probabilities of six specialty medical terms produced by this model with those obtained from models fine-tuned on the corresponding S-MedQA datasets. As shown in Table 6, there is a significant drop in token probabilities when fine-tuned on out-of-domain data, confirming that the change in token probabilities is brought by domain shifting rather than only addition training.
Table 6: Token probability increase by clinical specialty after fine-tuning.
Specialty Increase (%)
Cardio 10.5
Gastro 10.7
Infect 9.5
Neuro 8.4
Obstetrics 7.2
Pediatrics 7.5
@@
(Sample size) (80) (83) (102) (74) (88) (90)
Train Sets Llama-3.1 64.0 67.5 65.3 66.8 65.0 64.5 65.5
Cardio 69.3 76.5 73.5 76.2 70.5 70.9 72.8
Gastro 68.2 73.3 71.8 74.6 69.3 69.2 71.1
Infect 65.8 74.2 69.5 72.5 66.6 66.7 69.3
Neuro 69.9 76.4 72.3 76.9 69.5 71.7 72.8
Obstetrics 69.8 75.0 72.9 75.4 68.7 69.4 71.9
Pediatrics 66.9 75.0 72.1 75.9 68.7 70.2 71.4
Combined 72.9 79.5 75.4 77.8 70.8 74.1 75.1
Table 10: Cross-specialty accuracy matrix of Llama-3.1-8B-Instruct.

### v3

@@ Results.
Table 4 shows the performance of Llama-3.1-8B-Instruct fine-tuned independently on each specialty and tested on all six specialty test sets, as well as after fine-tuning on a combination of all six specialties. Are improvements truly indicative of knowledge acquisition or injection? Notably, for models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (69.9%) was achieved by the model trained on Neuro. If the improvements were due to knowledge injection, we would expect to see best-performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other LLMs in §A.5 and note that similar inconsistencies hold, whereas with different transfer patterns across specialties. This inconsistency suggests that the performance gains from fine-tuning primarily arise from factors other than knowledge injection, partially supporting Zhou et al. (2024)’s hypothesis in the clinical domain.
3.3 Term overlap analysis
Clinical questions from different specialties can differ substantially in terms of knowledge; for example, one cannot assume that neurology expertise can directly apply to cardiology. Here, we thus quantify the overlap of clinically relevant terms in train/test splits within and across specialties. We do this by extracting clinically relevant terms in each question of S-MedQA, by mapping these terms onto relevant clinical specialties within the scope of the top 6 specialties we selected for further experiments, and finally by quantifying the difference in clinical terminology across different specialties.
Figure 4: Heatmap showing the overlap of clinical terms within and across clinical specialties between train and test QA pairs in S-MedQA.
We map each term in a question to relevant clinical specialties using SNOMED-CT (Cornet and de Keizer, 2008) [^3] https://www.snomed.org/ and optionally the Human Phenotype Ontology (HPO; Castellanos et al., 2024). 1) We first manually identify the top-level disorder concepts in SNOMED-CT that correspond to each medical specialty, and use the SNOMED hierarchy to link specific diseases or concepts to this top-level category. E.g., we link any disorder under Disorder of cardiovascular system as part of cardiology. [^4] For all high-level SNOMED-CT concepts we use for each clinical specialty, please refer to § A.6. We found no simple mechanism to map findings, procedures, and observable entities concepts to disorders in SNOMED-CT directly. We thus first search findings, procedures, and observable entities in HPO, identify the related disorders, which we then use to map onto clinical specialties similarly with SNOMED-CT. 2) For each question, we then use SciSpacy Neumann et al. (2019) to perform entity linking of disorder, finding, procedure, and observable entity mentions in the question to SNOMED-CT. 3) Finally, we use the top-level concepts identified in (1) and the SNOMED-CT hierarchy to link each mention to a clinical specialty.
We exclude from this analysis general terms unrelated to specific specialties or shared by four or more of the six specialties.
Results.
Figure 4 illustrates the overlap percentages of specialty-specific terminology between training and test sets for the six specialties. The average overlap in train/test splits within the same specialty is 63.4%, indicating that training sets cover well clinical terminologies necessary for solving corresponding test set questions. In contrast, the average overlap across different specialties is only 32.8%, showing the domain-specific nature of medical terminology and the limited commonality between specialties. These results confirm that “knowledge leakage” during cross-specialty evaluations is minimal, as different specialties share limited common specialty-specific terminology.
3.4 Token probability
We hypothesize that the performance improvements in Table 4 are primarily due to domain shifting from general LLMs to the clinical domain, rather than the injection of new clinical knowledge. We demonstrate this point through a detailed analysis of token probability shifts for clinical terms before and after fine-tuning.
The impact of medical data fine-tuning.
To assess that, we analyze token probability changes for clinical terms in the questions from testsets linked to a single clinical specialty between a baseline model and the same model fine-tuned on data from each different specialty. Clinically relevant terms are extracted and mapped to specialties using the method described in §3.3, and each term’s probability is obtained by summing the probabilities of its constituent subword tokens.
In Fig. 5 we show the average log-probabilities of medical terms across specialties as predicted by Llama-3.1-8B-Instruct (base and fine-tuned) for each of the six specialties. Regardless of the specialty used for fine-tuning, we observe similarly increased token log-probabilities for clinical terms specific to the fine-tuning specialty, as well as for terms associated with other specialties.
We note that token log-probabilities for terms from different specialties differ in range, likely reflecting the pre-trained model’s existing knowledge distribution. This also seems to support our hypothesis that fine-tuning shifts the domain rather than injects new domain knowledge.
Are improvements due to additional training steps?
To rule out the possibility that the similarly increased probabilities in all clinical specialties are due simply to the additional training steps, we also fine-tune our baseline Llama-3.1-8B-Instruct on three social sciences subsets of MMLU (Hendrycks et al., 2021b, a)—public relations, security studies, and sociology—which are entirely unrelated to the medical domain. Again, we compare the average token log-probabilities of the six specialty medical terms produced by this model with those obtained from models fine-tuned on the corresponding S-MedQA datasets. When fine-tuning on MMLU out-of-domain datasets there is a significant drop in token probabilities compared to models trained on S-MedQA data (red bars in Figure 5). We thus confirm that this drop is caused by domain shifting rather than purely the effect of additional training. See Appendix A.7 for details of the token probabilities for models trained on S-MedQA and MMLU social sciences.
Figure 5: Negative log-probabilities for clinically relevant tokens between baseline Llama-3.1-8B-Instruct and the same model further fine-tuned on each specialty data. Each group represents tokens categorized into different clinical specialties. Each color means that the same model is further fine-tuned on each specialty data.
@@ 4 Discussion
The effectiveness of lightweight fine-tuning on strong base models has been widely recognized in NLP, largely following the pretraining paradigm introduced in the BERT era (Devlin et al., 2019). This strength is often attributed to domain shifting, especially for linguistically focused tasks such as sentiment analysis (Zhang et al., 2020), natural language inference (Koroteev, 2021), and semantic matching (Gao et al., 2021).
In knowledge-intensive domains such as clinical specialties, the role of fine-tuning remains underexplored. Clinical questions across specialties often pose substantial knowledge barriers, and the capabilities of current LLMs on specialty-specific tasks are still unclear. Moreover, it remains uncertain whether these tasks benefit more from explicit knowledge injection or from effective knowledge transfer within the base model. This paper benchmarks these real-world questions in a fine-grained manner and provides preliminary evidence to explain the role of fine-tuning in the clinical domain.
Moreover, current LLMs are trained on massive datasets and raise concerns about benchmark contamination. Our experiments span a wide range of models, including open-source ones such as OLMo (Groeneveld et al., 2024), which is trained without clinical data. Similar findings on these models strengthen the reliability of our results.
5 Related Work
Medical QA datasets.
Medical QA datasets have played a pivotal role in benchmarking and advancing the capabilities of LLMs in the healthcare domain. Early datasets such as PubMedQA (Jin et al., 2019) and MedQA (Jin et al., 2021) introduced multiple-choice QA tasks derived from medical board exams, focusing on assessing clinical reasoning and factual knowledge. MedMCQA (Pal et al., 2022) later expanded the scope with a larger dataset covering a wide range of medical subjects. These datasets have been instrumental in developing models like BioGPT (Luo et al., 2022) and MedPaLM (Singhal et al., 2025), which leverage domain-specific pretraining to achieve human-like performance on medical tasks.
However, these datasets lack annotations for clinical specialties, limiting their applicability in scenarios where specialty-specific knowledge is crucial. Recent efforts, such as BioMistral (Labrak et al., 2024), have focused on pretraining biomedical LLMs on medical corpora but have not addressed the need for fine-grained specialty-level evaluation. S-MedQA fills this gap by introducing per-specialty annotations and enabling the study of knowledge transfer across specialties.
Knowledge injection vs. Domain shifting.
The effectiveness of fine-tuning on task-specific datasets has been extensively studied in NLP (Devlin et al., 2019; Mayfield and Black, 2020; Limsopatham, 2021). Broadly speaking, this effectiveness is often understood to either introduce new knowledge within the external corpus into pre-trained models (Fu et al., 2023; Mecklenburg et al., 2024), especially in knowledge-intensive tasks, or to adapt a model’s general capabilities to more specialized domains (Limsopatham, 2021; Garrido-Merchán et al., 2023; Wu et al., 2024), highlighting effective knowledge sharing across tasks.
Particularly, recent studies (Zhou et al., 2024) in LLMs suggest that fine-tuning primarily serves to shift the model’s focus toward a specific domain, leveraging knowledge already encoded during pretraining, rather than injecting new knowledge into the model. Zhou et al. (2024) introduced the hypothesis that alignment tuning—e.g., supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF)—primarily teaches large language models (LLMs) to select a sub-distribution of response styles and formats. This hypothesis was supported by their study on the LIMA model, where fine-tuning with as few as 1,000 examples yielded strong alignment. This study demonstrated that the success of alignment tuning lies in its ability to adjust stylistic tokens, transitional phrases, and disclaimers, while the majority of the knowledge required to answer user queries remains unchanged from the base model. Our work is a first step towards investigating the role of fine-tuning in knowledge-intensive and finely segmented domains, since we specifically address the medical domain with different clinical specialties.
@@ 6 Conclusions
In this paper, we introduce S-MedQA, the first English medical QA dataset annotated for 15 distinct clinical specialties including single- and multi-specialty annotations for cross-disciplinary questions. We investigate different questions using S-MedQA and find that questions from different clinical specialties have significant differences in terms of clinical terminology. Despite that, we also find that the best QA performance is almost always observed for LLMs fine-tuned on unrelated clinical specialties. Moreover, we also show that fine-tuning the same base LLM on high-quality non-medical data leads to decreased token probabilities for clinically relevant terms. Our results suggest that Zhou et al. (2024)’s hypothesis apply to the medical domain and that improvements can be primarily attributed to domain shifting rather than knowledge injection. However, the precise impact of different types of QA data (e.g., complexity or difficulty of the QA pair) remains unclear. Finally, we recommend further research to investigate the role of fine-tuning in the medical domain.
@@
In the following, we discuss some limitations of our work.
Experiments on knowledge transfer patterns
Although we conduct experiments investigating how knowledge transfers within and across clinical specialties, we limited our experiments to the medical domain. That means that validating to what extent Zhou et al. (2024)’s hypothesis generalises to other knowledge-intensive domains is not something we address in this work. Further research is needed to investigate the role of fine-tuning and instruction data in other domains.
Clinical specialties
We have included 15 clinical specialties in our dataset due to their availability in the source datasets we used. This means that possibly relevant clinical specialties have been left out of our dataset (i.e., see the long tail in Figure 2). Future work should aim to include QA pairs from more varied and possibly rarer clinical specialties (e.g., speech language pathology;  Kim et al. 2024).
Language and healthcare system
Our dataset draws on examples in English extracted mostly from US-centric datasets. We believe that important future work lies in extending QA datasets to other languages and healthcare systems.
@@ A.2 Prompts used for specialty classification
A.3 Prompts used for LLM tuning and inferring
A. Atenolol
B. Diltiazem
C. Propafenone
D. Digoxin
Answer: [/INST] D. Digoxin
@@ D. Digoxin
A.5 Additional cross-specialty evaluation results
@@
In Tables 56,  7,  4,  9,  10,  11 we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, Mistral-7b-instruct-v0.1, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Bio-Medical-LLaMA-3-8B, and OLMo-2-1124-7B-Instruct in addition to our main results (in §3.2). Here we also observe that the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty (i.e., we observe this in $\sim$70% of the cases).
@@
Mistral-v0.2 52.0 45.9 48.2 37.0 52.9 43.5 46.9
Train Sets Cardio 51.8 54.6 44.3 47.6 51.5 44.6 49.4
Gastro 54.9 54.1 38.9 43.5 51.7 46.4 48.8
Infect 55.4 53.9 43.0 47.8 54.3 43.5 50.1
Neuro 54.4 54.8 41.9 45.7 56.0 49.0 50.8
Obstetrics 54.4 52.4 45.7 41.9 51.5 44.6 48.8
Pediatrics 50.8 53.0 42.4 35.9 48.1 46.1 46.5
Combined 53.7 53.8 42.7 43.7 52.2 45.7 49.1
Table 8: Cross-specialty accuracy matrix of Mistral-v0.1
@@
A.6 High-level Concepts in SNOMED-CT
The high-level SNOMED-CT disorder concepts we use when mapping questions to clinical specialities are: cardiology (Disorder of cardiovascular system (SCTID: 49601007)), gastroenterology (Disorder of digestive system (SCTID: 53619000)), infectious diseases (Infectious disease (SCTID: 40733004)), obstetrics (Disorder of female reproductive system (SCTID: 363124003), Disorder of fetus and/or mother during labor (SCTID: 1269083008)), neurology (Disorder of nervous system(SCTID: 118940003)), pediatrics (Behavioral and emotional disorder with onset in childhood (SCTID: 231538003), Disorder of fetus and/or newborn (SCTID: 414025005), Developmental disorder (SCTID: 5294002)).
A.7 Token probability change
Table 12 shows the token negative log-likelihoods (the higher the better) for terms from each clinical specialty before and after fine-tuning on the six S-MedQA specialty data and MMLU social sciences data. We note that, apart from a few exceptions, training on S-MedQA leads to consistently improved token probabilities for clinically relevant terms, whereas training on MMLU makes clinical relevant terms less likely.
Models Test Sets ($\uparrow$ is better)
Cardio Gastro Infect Neuro Obstetrics Pediatrics
Mistralv0.2-Base -4.11 -5.05 -4.99 -5.59 -4.49 -5.80
Cardio -4.06 -4.94 -4.86 -5.56 -4.35 -5.60
Gastro -4.12 -5.01 -4.91 -5.46 -4.27 -5.65
Infect -3.95 -4.84 -4.96 -5.55 -4.41 -5.59
Neuro -4.07 -5.05 -4.88 -5.39 -4.37 -5.61
Obstetrics -4.14 -5.11 -4.87 -5.50 -4.41 -5.66
Pediatrics -4.08 -5.05 -4.93 -5.53 -4.46 -5.67
MMLU -4.62 -5.34 -5.27 -5.83 -4.78 -6.01
Table 12: Average token negative log-likelihood for terms from each clinical specialty before and after fine-tuning on the six S-MedQA specialty data and MMLU social sciences data. Best viewed in colour.